# Tackling Decision Processes with Non-Cumulative Objectives using Reinforcement Learning

### v1

@@ 3 Experiments
From an implementation perspective, our scheme of mapping NCMDPs to MDPs requires minimal effort, since we can treat both the NCMDP and the used MDP solver as black boxes by simply adding a layer between them, as shown in Figure 1. In addition, our treatment facilitates online learning and does not require any computationally expensive preprocessing of the NCMDP. This opens the door for researchers with specific domain knowledge, who are not necessarily experts in reinforcement learning, to use standard libraries such as stable-baselines3 [Raffin et al. (2021)] to solve their non-cumulative problems. Conversely, it allows reinforcement learning experts to quickly tackle existing environments using our method.
In the following, we focus on reinforcement learning (RL). We provide details on hyperparameters, compute resources, and training curves for all experiments in Appendix A.
@@ 3.1 Classical Control
Figure 3: Trade-off between average cumulative reward in the Lunar lander environment (inset) and maximum speed increase for the non-cumulative MAXVELPPO and the cumulative SUMVELPPO algorithm.
As a first use case of our method, we train a reinforcement learning agent in the Lunar lander environment of the gymnasium [Towers et al. (2023)] library. The agent controls a spacecraft with four discrete actions corresponding to different engines while being pushed by a stochastic wind. Immediate positive rewards $r_{t}$ are given for landing the spacecraft safely with small negative rewards given for using the engines. A realistic goal when landing a spacecraft is to not let the spacecraft get too fast, e.g. to avoid excessive frictional heating. Therefore, we define an NCMDP where the agent is penalized for its maximum speed during a trajectory, i.e. we try to maximize
@@
where $v_{t}$ is the speed of the agent at time $t$ and $c$ defines a trade-off between minimizing the maximum speed and the other goals of the agent. We train RL agents using Proximal Policy Optimization (PPO) [Schulman et al. (2017)] on the MDP constructed from the NCMDP as described above (MAXVELPPO). To show the trade-off between the sum of the agents’ original cumulative rewards $\sum_{t=0}^{T-1}r_{t}$ and their maximum speed increase during a trajectory, we train agents for different $c$. Each value of $c$ corresponds to a marker in Figure 3. We compare our results to an RL agent with a similar but cumulative objective maximizing $\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}(r_{t}-cv_{t}^{2})\right]$ (SUMVELPPO). As shown in Figure 3, the non-cumulative MAXVELPPO agent is consistently able to find a better trade-off than the cumulative SUMVELPPO agent. All experiments were performed with 5 agents using different seeds. We plot average results and standard deviation as error bars for both the cumulative reward and the speed increase (mostly hidden behind markers).
@@ 3.2 Portfolio Optimization with Sharpe Ratio as Objective
Figure 4: Portfolio optimization. Sharpe ratio during 10 different training periods for algorithms maximizing the cumulative differential Sharpe ratio (DIFFSHARPE) and the non-cumulative exact Sharpe ratio (SHARPE).
Table 2: Average Sharpe ratios.
DIFFSHARPE SHARPE
Training $10.51\pm 1.65$ $13.33\pm 1.14$
Evaluation $1.43\pm 0.18$ $1.35\pm 0.15$
Test $0.53\pm 0.26$ $0.51\pm 0.09$
Next, we consider the task of portfolio optimization where an agent decides how to best invest its assets across different possibilities. A common measure for a successful investment strategy is its Sharpe ratio [Sharpe (1966)]
$\frac{\mathrm{MEAN}(\tilde{r}_{0},\dots,\tilde{r}_{T-1})}{\mathrm{STD}(\tilde{r}_{0},\dots,\tilde{r}_{T-1})},$ (7)
where $\tilde{r}_{t}=(P_{t+1}-P_{t})/P_{t}$ are the simple returns and $P_{t}$ is the portfolio value at time $t$. By dividing through the standard deviation of the simple returns, the agent is discouraged from risky strategies with high volatility. As the Sharpe ratio is non-cumulative, reinforcement learning strategies so far needed to fall back on the approximate differential Sharpe ratio as a reward [Moody et al. (1998); Moody & Saffell (2001)]. However, using the methods developed in this paper we can directly maximize the exact Sharpe ratio. We perform experiments on an environment as described by Sood et al. (2023) where an agent trades on the 11 different S&P500 sector indices between 2006 and 2021. The states contain a history of returns of each index and different volatility measures. Actions are the continuous relative portfolio allocations for each day. The experiment described by Sood et al. (2023) consists of training 5 agents with different seeds over a 5-year period, periodically evaluating their performance on the following year, and testing the best-performing agent in the year after that. Then, the time period is shifted by one year into the future resulting in a total of 10 time periods. We re-implement this experiment by training agents using PPO with the cumulative differential Sharpe ratio (DIFFSHARPE) or the exact Sharpe ratio (SHARPE) as their objective. As depicted in Figure 4, the SHARPE algorithm significantly outperforms the DIFFSHARPE algorithm in the training years. However, as shown in Table 2, on the evaluation and test years there is no significant difference indicating over-fitting of the agents’ policies to the years they are trained on. Nonetheless, we expect training on the exact SHARPE ratio to give consistently better results if the problem of over-fitting is solved, e.g. by using realistic stock-market simulators which are for example being developed using Generative Adversarial Networks [Li et al. (2020)]. The experiments were performed with 5 sets of 5 seeds each and the standard deviation of different sets of seeds is reported.
@@ 3.3 Discrete Optimization Problems
Next, we consider a large class of applications where RL is commonly used: Problems where the agent iteratively transforms a state by its actions to find a state with a lower associated cost. These problems are common in scientific applications such as physics, e.g. to reduce the length of quantum logic circuits [Fösel et al. (2021)], or chemistry, e.g. for molecular discovery [Zhou et al. (2019)]. Another prominent example is the discovery of new algorithms [Mankowitz et al. (2023)]. Intuitively, these problems can be understood as searching for the state with the lowest cost within an equivalence class defined by all states that can be reached from the start state by the agent’s actions.
@@
We conjecture that maximizing Equation 8 will yield better results than maximizing $\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}\tilde{r}_{t}\right]$ due to the following reasons:
@@
Peak environment
To facilitate an in-depth analysis, we first consider a toy environment with the cost function depicted in the inset of Figure 5 (a). The cost function was chosen to be simple while still requiring intermittently cost-increasing actions of the optimal policy. Each trajectory lasts 10 steps and the agent’s actions are stepping to the left, right, or doing nothing. To minimize the number of hyperparameters, we use the REINFORCE algorithm [Williams (1992)] with a tabular policy. We compare agents trained with the non-cumulative objective Equation 8 (MAXREINFORCE) with agents that maximize the cumulative rewards (REINFORCE). As shown in Figure 5 (a), the MAXREINFORCE agent trains much faster. Two possible sources of this speed-up are a different, more advantageous direction of the gradient updates and a reduced variance when estimating these gradients. To investigate which is the case, we periodically stop training and run $n=1000$ trajectories with a fixed policy. We then compute the empirical variance of the gradient update inspired by Kaledin et al. (2022) as
$\mathrm{VAR}=\frac{1}{n}\sum_{i=0}^{n-1}\left\lVert\vec{g}_{i}\right\rVert_{2}^{2}-\left\lVert\frac{1}{n}\sum_{i=0}^{n-1}\vec{g}_{i}\right\rVert_{2}^{2},$ (9)
where $\vec{g}_{i}$ is the gradient from a single trajectory derived either by the MAXREINFORCE or the REINFORCE algorithm. As this variance scales with the squared magnitude of the average gradients, we normalize it by this value to ensure a fair comparison between the two algorithms. In Figure 5 (b), we show that in the initial phase of training, the MAXREINFORCE algorithm significantly reduces variance. To compare the direction of the gradients we use the same data to compute the normalized average gradients of both algorithms and show their dot product in Figure 5 (c). We find that the gradients are correlated (i.e. the dot product is bigger than zero) but not the same. Therefore, we conclude that the training speed-up is derived both from lower variance and a better true gradient direction. All results reported are averaged over 10 seeds with standard deviations plotted shaded.
@@
Figure 6: ZX environment. (a) Cost improvement during the trajectory against total steps taken in the environment for the cumulative REINFORCE agent and the MAXREINFORCE agent maximizing Equation 8. Left inset: Zoom-in to the initial training phase. Right inset: Zoom-in to the final training phase. (b) Entropy of the agents’ policies against total steps taken in the environment. Inset: A typical ZX-diagram.
ZX-diagrams
As a real-world use case, we consider the simplification of ZX-diagrams, which are graph representations of quantum processes [Coecke & Kissinger (2017)] with applications e.g. in the compilation of quantum programs [Duncan et al. (2020); Riu et al. (2023)]. An example of a typical ZX-diagram is shown in the inset of Figure 6 (b). We consider the environment described by Nägele & Marquardt (2023), where the cost function of a diagram is given by its node number, the start states are randomly sampled ZX-diagrams, and the actions are a set of local graph transformations. In total, there are 6 actions per node and 6 actions per edge in the diagram. This is a challenging reinforcement learning task that requires the use of graph neural networks to accommodate the changing size of the state and action space. We use PPO to train agents to maximize either Equation 8 (MAXPPO) or the cumulative reward (PPO) with a trajectory length of 20 steps. As shown in Figure 6 (a) the MAXPPO agent initially trains faster than the PPO agent (left inset). This is likely due to the reduced variance and different gradient direction as described above. The PPO agent then shortly catches up, but ultimately requires about twice as many training steps to reach optimal performance as the MAXPPO agent (right inset). We argue that this is because the MAXPPO agent is better at exploring and therefore learning difficult optimization strategies (reason 3 above). This is captured by the entropy of the MAXPPO agent’s policy staying much higher than the entropy of the PPO agent, as shown in Figure 6 (b). All reported results are averaged over 5 seeds with standard deviations plotted shaded.
Quantum error correction
@@
Next, we focus on optimization problems where the starting state is always the same and the optimal stopping point is known. The specific tasks we consider are the search for quantum programs to either find new quantum error correction codes or to prepare logical states of a given quantum error correction code - both critical for the eventual realization of quantum computation [Terhal (2015)]. In both cases, the agent iteratively adds elementary quantum logic gates until the program delivers the desired result. The agents’ policies are encoded by standard multilayer perceptrons. We train agents on five different sized problems of both tasks. For details on the used environment, see Zen et al. (2024) and Olle et al. (2023). To obtain a single performance measure encompassing both final performance and training speed, we continuously evaluate the mean cost improvement of the agents and average it over the training process as suggested by Andrychowicz et al. (2020b). Intuitively, this measure can be understood as the area under the agent’s training curve. In Table 3, we report the quotient of this performance measure of the MAXPPO algorithm and the PPO algorithm for both the best of 10 and the mean of 10 trained agents. The different tasks are denoted by three integers, as customary in the quantum error correction community. We find that the MAXPPO algorithm performs similarly to PPO in quantum error correction code discovery (subscript e), and significantly better in logical state preparation (subscript s). We argue that the large degeneracy present in the solution space of the code discovery task diverts the more exploratory MAXPPO, lowering its performance to the level of PPO.
Table 3: Results for quantum error correction environments.
@@
$[[5,1,3]]_{\mathrm{e}}$ $\mathbf{1.0004}$ $\mathbf{1.0001\pm 0.0008}$
$[[7,1,3]]_{\mathrm{e}}$ $0.9993$ $0.9994\pm 0.0003$
$[[9,1,3]]_{\mathrm{e}}$ $0.9994$ $0.9994\pm 0.0003$
$[[10,2,4]]_{\mathrm{e}}$ $\mathbf{1.0013}$ $\mathbf{1.0023\pm 0.0017}$
$[[12,2,4]]_{\mathrm{e}}$ $0.9993$ $0.9991\pm 0.0007$
$[[5,1,3]]_{\mathrm{s}}$ $\mathbf{1.0419}$ $\mathbf{1.0666\pm 0.0407}$
$[[7,1,3]]_{\mathrm{s}}$ $\mathbf{1.1824}$ $\mathbf{1.2090\pm 0.0623}$
$[[9,1,3]]_{\mathrm{s}}$ $\mathbf{1.0060}$ $\mathbf{1.0905\pm 0.1079}$
$[[15,1,3]]_{\mathrm{s}}$ $\mathbf{1.1382}$ $\mathbf{1.0309\pm 0.0806}$
$[[17,1,5]]_{\mathrm{s}}$ $\mathbf{1.4748}$ $\mathbf{1.6218\pm 0.2144}$
Limitation
When applying MAXPPO to discrete optimization problems with long trajectories in the many hundreds of steps, we empirically observed an initially slow learning speed. This could be due to the agent initially mostly increasing cost and therefore receiving zero reward for almost the entire trajectory. A possible solution could be to dynamically adjust the trajectory length in these problems going from shorter to longer trajectories during the training process.
@@ 4 Related Work
A special case of NCMDPs is first considered by Quah & Quek (2006), who adapt Q-learning to the $\max$ objective by redefining the temporal difference error of their learning algorithms and demonstrate their algorithms on an optimal stopping problem. However, they do not adapt state space and do not provide theoretical convergence guarantees. Gottipati et al. (2020) rediscover the same algorithm, apply it to molecule generation, and provide convergence guarantees for their method, while Eyckerman et al. (2022) consider the same algorithm for application placement in fog environments. Cui & Yu (2023) finally show a shortcoming of the method used in the above papers: It is guaranteed to converge to the optimal policy only in deterministic settings. Additionally, they provide convergence guarantees of Q-learning in deterministic environments for a larger class of functions $f$, focusing on the $\min$ function for network routing applications. Independently, the $\max$ function is also used in the field of safety reinforcement learning in deterministic environments both for Q-learning [Fisac et al. (2015, 2019); Hsu et al. (2021)] and policy-based reinforcement learning [Yu et al. (2022)]. Moflic & Paler (2023) use a reward function with parameters that depend on the past rewards of the trajectory to tackle quantum circuit optimization problems that require large intermittent negative rewards, albeit without providing theoretical convergence guarantees. Additionally, Wang et al. (2020) investigate planning problems with non-cumulative objectives within deterministic settings. They provide provably efficient planning algorithms for a large class of functions $f$ by discretizing rewards and appending them to the states. Recently, Veviurko et al. (2024) showed how to use the $\max$ objective also in probabilistic settings by augmenting state space and provide convergence guarantees for both Q-learning and policy-based methods. They then show experiments with their algorithm yielding improvements in MDPs with shaped rewards. However, they do not redefine the rewards, which requires adaptation of the implementation of their MDP solvers. Our method described above reduces to an effectively equivalent algorithm to theirs in the special case of the $\max$ objective.

### v3

@@ 3 Experiments
Figure 3: Average return in the lunar lander environment (inset) against the velocity penalty coefficient $c$ defined in Equation 7 for the non-cumulative MAXVELPPO and the cumulative FINALPPO algorithm.
There are three kinds of experiments presented in this manuscript:
  1. 1.
    We show that our method, in conjunction with dynamic programming, can find the optimal policy in a proof-of-principle stochastic NCMDPs where previous methods fail (see previous section, Figure 2).
  2. 2.
    In the supplemental Appendix C, we compare our method with the only previous approach for solving NCMDPs with objectives other than the $\max$ function, which was introduced by Cui and Yu (2023), and is only applicable with more restrictive objectives $f$, in deterministic environments, and in conjunction with Q-learning methods. Our experiments indicate that even in this restricted setting, our more general method may outperform the previous method in terms of final performance.
  3. 3.
    In the following sections, we show applications in real-world tasks where previous methods used for solving special cases of NCMDPs cannot be applied. In lack of a general method for solving NCMDPs, state-of-the-art approaches to these real-world problems have so far relied on standard MDPs and approximate solutions which we use as a baseline to compare our method to.
If not otherwise specified, shaded regions in plots and uncertainties in tables denote a $95\%$ bootstrap confidence interval calculated over multiple seeds of the experiment. We provide details on hyperparameters, compute resources, and training curves for all experiments in the supplemental Appendix D.
@@ 3.1 Classical Control
As a first use case of our method, we train a reinforcement learning agent in the lunar lander environment of the gymnasium (Towers et al., 2023) library. The agent controls a spacecraft with four discrete actions corresponding to different engines while being pushed by a stochastic wind. Immediate positive rewards $r_{t}$ are given for landing the spacecraft safely with small negative rewards given for using the engines. A realistic goal when landing a spacecraft is to not let the spacecraft get too fast, e.g. to avoid excessive frictional heating. Therefore, we define an objective where the agent is penalized for its maximum speed during a trajectory, i.e. we try to maximize
@@
where $v_{t}$ is the speed of the agent at time $t$ and $c$ defines a trade-off between minimizing the maximum speed and the other goals of the agent. We train RL agents using Proximal Policy Optimization (PPO) (Schulman et al., 2017) on the MDP constructed from the NCMDP as described above (MAXVELPPO) for different values of $c$. We compare our results to an RL agent where the velocity penalty $-c\max{(v_{0},\dots,v_{T-1})}$ is added to the final reward resulting in a cumulative objective (FINALPPO). As shown in Figure 3, the non-cumulative MAXVELPPO agent can find a better trade-off than the cumulative FINALPPO agent.
The reward function of the FINALPPO agent is non-Markovian. Therefore, we also tested adding the extra state information $h_{t}$, i.e. the past maximum velocity, to the FINALPPO agent’s observations to restore Markovianity. However, this did not result in a performance gain (see Figure A3). All experiments were performed with 5 agents using different seeds.
@@ 3.2 Portfolio Optimization with Sharpe Ratio as Objective
Figure 4: Portfolio optimization. Sharpe ratio during 10 different training periods for algorithms maximizing either the cumulative differential Sharpe ratio (DIFFSHARPE) or the exact Sharpe ratio by using our non-cumulative method (SHARPE) or by giving the Sharpe ratio as a single final reward (FINALSHARPE).
Next, we consider the task of portfolio optimization where an agent decides how to best invest its assets across different possibilities. A common measure for a successful investment strategy is its Sharpe ratio (Sharpe, 1966)
$\frac{\mathrm{MEAN}(\tilde{r}_{0},\dots,\tilde{r}_{T-1})}{\mathrm{STD}(\tilde{r}_{0},\dots,\tilde{r}_{T-1})},$
where $\tilde{r}_{t}=(P_{t+1}-P_{t})/P_{t}$ are the simple returns and $P_{t}$ is the portfolio value at time $t$. By dividing through the standard deviation of the simple returns, the agent is discouraged from risky strategies with high volatility. As the Sharpe ratio is non-cumulative, reinforcement learning strategies so far needed to fall back on the approximate differential Sharpe ratio as a reward (Moody et al., 1998; Moody and Saffell, 2001). However, using the methods developed in this paper we can directly maximize the exact Sharpe ratio. We perform experiments on an environment as described by Sood et al. (2023) where an agent trades on the 11 different S&P500 sector indices between 2006 and 2021. The states contain a history of returns of each index and different volatility measures. Actions are the continuous relative portfolio allocations for each day. The experiment described by Sood et al. (2023) consists of training 5 agents with different seeds on 10 overlapping periods spanning 5 years each. We re-implement this experiment by training agents using PPO with the cumulative differential Sharpe ratio (DIFFSHARPE) or the exact Sharpe ratio as their objective. For the agents maximizing the exact Sharpe ratio, we either phrase the problem as an NCMDP and use our method described above (SHARPE), or we provide a single reward at the end of the trajectory (FINALSHARPE). As depicted in Figure 4, the SHARPE algorithm significantly outperforms the other algorithms. As in Section 3.1, restoring Markovianity in the FINALSHARPE algorithm by adding $h_{t}$ to the states did not improve performance.
Limitation: After training, all of the considered algorithms do not generalize well to new, unseen periods indicating over-fitting of the agents’ policies to the years they are trained on. Therefore, the experiments presented here should be understood as showing improved performance of the SHARPE agent only in the environment it was trained in. This limitation could potentially be offset by generating more data to train on, e.g. by using realistic stock-market simulators which are for example being developed using Generative Adversarial Networks (Li et al., 2020).
@@ 3.3 Discrete Optimization Problems
Next, we consider a large class of applications where RL is commonly used: Problems where the agent iteratively transforms a state by its actions to find a state with a lower associated cost. These problems are common in scientific applications such as physics, e.g. to reduce the length of quantum logic circuits (Fösel et al., 2021), or chemistry, e.g. for molecular discovery (Zhou et al., 2019). Another prominent example is the discovery of new algorithms (Mankowitz et al., 2023). Intuitively, these problems can be understood as searching for the state with the lowest cost within an equivalence class defined by all states that can be reached from the start state by the agent’s actions.
@@
In the deterministic settings we study in the following experiments, the ideal policy for maximizing the cumulative reward $\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}\tilde{r}_{t}\right]$ maximizes also Equation 8. We conjecture that directly maximizing Equation 8 by formulating the problem as an NCMDP will yield better results due to the following reasons:
@@
3.3.1 Peak environment
To facilitate an in-depth analysis, we first consider a toy environment with the cost function depicted in the inset of Figure 5 (a). The cost function was chosen to be simple while still requiring intermittently cost-increasing actions of the optimal policy. Each trajectory lasts 10 steps and the agent’s actions are stepping to the left, right, or doing nothing. To minimize the number of hyperparameters, we use the REINFORCE algorithm (Williams, 1992) with a tabular policy. We compare agents trained with the non-cumulative objective Equation 8 (MAXREINFORCE) with agents that maximize the cumulative rewards (REINFORCE). As shown in Figure 5 (a), the MAXREINFORCE agent trains much faster. Two possible sources of this speed-up are a different, more advantageous direction of the gradient updates and a reduced variance when estimating these gradients. To investigate which is the case, we periodically stop training and run $n=1000$ trajectories with a fixed policy. We then compute the empirical variance of the gradient update inspired by Kaledin et al. (2022) as
$\mathrm{VAR}=\frac{1}{n}\sum_{i=0}^{n-1}\left\lVert\vec{g}_{i}\right\rVert_{2}^{2}-\left\lVert\frac{1}{n}\sum_{i=0}^{n-1}\vec{g}_{i}\right\rVert_{2}^{2},$
where $\vec{g}_{i}$ is the gradient from a single trajectory derived either by the MAXREINFORCE or the REINFORCE algorithm. As this variance scales with the squared magnitude of the average gradients, we normalize it by this value to ensure a fair comparison between the two algorithms. In Figure 5 (b), we show that in the initial phase of training, the MAXREINFORCE algorithm significantly reduces variance. To compare the direction of the gradients we use the same data to compute the normalized average gradients of both algorithms and show their dot product in Figure 5 (c). We find that the gradients are correlated (i.e. the dot product is bigger than zero) but not the same. Therefore, we conclude that the training speed-up is derived both from lower variance and a better true gradient direction. All results reported are averaged over 10 seeds.
3.3.2 Quantum error correction
@@
As a first real-world use case, we focus on optimization problems where the initial state is always the same and the optimal stopping point is known. The specific task we consider is the search for quantum programs to prepare logical states of a given quantum error correction code - critical for the eventual realization of quantum computation (Terhal, 2015). The agent iteratively adds elementary quantum logic gates until the program delivers the desired result. The agents’ policies are encoded by standard multilayer perceptrons and they are trained on five different sized problems. For details on the used environment, see Zen et al. (2024). We use PPO to train agents to maximize either Equation 8 (MAXPPO) or the cumulative reward (PPO). To obtain a single performance measure encompassing both final performance and training speed, we continuously evaluate the mean cost improvement of the agents and average it over the training process as suggested by Andrychowicz et al. (2020b). Intuitively, this measure can be understood as the area under the agent’s training curve. In Table 2, we report the quotient of this performance measure of the MAXPPO algorithm and the PPO algorithm for both the best of 10 and the mean of 10 trained agents. The different tasks are denoted by three integers, as customary in the quantum error correction community. We find that the MAXPPO algorithm performs significantly better.
We also trained MAXPPO for a second quantum circuit discovery task [for details see Olle et al. (2024)]. In this case, PPO and MAXPPO performed equally well. We argue that the large degeneracy present in the solution space of this second discovery task diverts the more exploratory MAXPPO, lowering its performance to the level of PPO.
@@
$[[5,1,3]]$ $1.04$ $1.07\,(1.04,1.09)$
$[[7,1,3]]$ $1.18$ $1.21\,(1.17,1.25)$
$[[9,1,3]]$ $1.01$ $1.09\,(1.04,1.17)$
$[[15,1,3]]$ $1.14$ $1.03\,(0.98,1.08)$
$[[17,1,5]]$ $1.47$ $1.62\,(1.52,1.78)$
Table 2: Results for quantum error correction environments.
3.3.3 ZX-diagrams
As another real-world discrete optimization problem, we consider the simplification of ZX-diagrams, which are graph representations of quantum processes (Coecke and Kissinger, 2017) with applications e.g. in the compilation of quantum programs (Duncan et al., 2020; Riu et al., 2023). An example of a typical ZX-diagram is shown in the inset of Figure 6 (b). We consider the environment described by Nägele and Marquardt (2024), where the cost function of a diagram is given by its node number, the start states are randomly sampled ZX-diagrams, and the actions are a set of local graph transformations. In total, there are 6 actions per node and 6 actions per edge in the diagram. This is a challenging reinforcement learning task that requires the use of graph neural networks to accommodate the changing size of the state and action space. As shown in Figure 6 (a) the MAXPPO agent initially trains faster than the PPO agent (left inset). This is likely due to the reduced variance and different gradient direction as described above. The PPO agent then shortly catches up, but ultimately requires about twice as many training steps to reach optimal performance as the MAXPPO agent (right inset). We argue that this is because the MAXPPO agent is better at exploring and therefore learning difficult optimization strategies (reason 3 in Section 3.3). This is captured by the entropy of the MAXPPO agent’s policy staying much higher than the entropy of the PPO agent, as shown in Figure 6 (b). The reported results are averaged over 5 seeds.
Figure 6: ZX environment. (a) Cost improvement during the trajectory against total steps taken in the environment for the cumulative PPO agent and the MAXPPO agent maximizing Equation 8. Left inset: Zoom-in to the initial training phase. Right inset: Zoom-in to the final training phase. (b) Entropy of the agents’ policies against total steps taken in the environment. Inset: A typical ZX-diagram.
Limitation: When applying MAXPPO to discrete optimization problems with long trajectories in the many hundreds of steps, we empirically observed an initially slow learning speed. This could be due to the agent initially mostly increasing cost and therefore receiving zero reward for almost the entire trajectory. A possible solution could be to dynamically adjust the trajectory length in these problems going from shorter to longer trajectories during the training process.
@@ 4 Related Work
A special case of NCMDPs is first considered by Quah and Quek (2006), who adapt Q-learning to the $\max$ objective by redefining the temporal difference error of their learning algorithms and demonstrate their algorithms on an optimal stopping problem. However, they do not adapt state space and do not provide theoretical convergence guarantees. Gottipati et al. (2020) rediscover the same algorithm, apply it to molecule generation, and provide convergence guarantees for their method, while Eyckerman et al. (2022) consider the same algorithm for application placement in fog environments. Cui and Yu (2023) finally show a shortcoming of the method used in the above papers: It is guaranteed to converge to the optimal policy only in deterministic settings. Additionally, they provide convergence guarantees of Q-learning in deterministic environments for a larger class of functions $f$, focusing on the $\min$ function for network routing applications. In Section B.1 we show that the class of objectives $f$ considered by Cui and Yu (2023) is a subclass of all objectives that lead to constant size extra state information $h_{t}$ using our method. Independently, the $\max$ function is also used in the field of safety reinforcement learning in deterministic environments both for Q-learning (Fisac et al., 2015, 2019; Hsu et al., 2021) and policy-based reinforcement learning (Yu et al., 2022). Moflic and Paler (2023) use a reward function with parameters that depend on the past rewards of the trajectory to tackle quantum circuit optimization problems that require large intermittent negative rewards, albeit without providing theoretical convergence guarantees. Additionally, Wang et al. (2020) investigate planning problems with non-cumulative objectives within deterministic settings. They provide provably efficient planning algorithms for a large class of functions $f$ by discretizing rewards and appending them to the states. Recently, Veviurko et al. (2024) showed how to use the $\max$ objective also in probabilistic settings by augmenting state space and provide convergence guarantees for both Q-learning and policy-based methods. They then show experiments with their algorithm yielding improvements in MDPs with shaped rewards. However, they do not redefine the rewards, which requires adaptation of the implementation of their MDP solvers. Our method described above reduces to an effectively equivalent algorithm to theirs in the special case of the $\max$ objective.